OpenStudy (sonyalee77):

Determine which ordered pair is NOT a solution of y=x+10. A.(-6,4) B.(-1,9) C.(-8,3) D.(4,14)

ganeshie8 (ganeshie8):

you will have to try each of the given option and see which one satisfies

ganeshie8 (ganeshie8):

or NOT satisfies

OpenStudy (anonymous):

It would be C.

ganeshie8 (ganeshie8):

for ex, lets option A : (-6, 4) y = x + 10 plugin x = -6, and see if you get 4 y = -6 + 10 =4 so (-6, 4) satisfies the given equaiton

ganeshie8 (ganeshie8):

similarly test other options, until u found the one which doesnt satisfy

OpenStudy (sonyalee77):

thanx :D

ganeshie8 (ganeshie8):

np.. u wlc :)